OpenStudy (anonymous):

Can someone help me with this one Find the area of the following ellipse. (Round answer to the nearest tenth). 2a = 10 cm; 2b = 20 cm. A = _____cm²

OpenStudy (anonymous):

The area of an ellipse is given by A = pi * a * b. So we would have: A = pi( 5 )( 10 ) = 157.0796327 To the nearest tenth, you have: \[A = 157.1cm^{2}\] I'm assuming that 2a is twice the length of half of the semi-major axis and 2b twice the length of half the semi-minor axis. Follow this link to get more information: http://www.math.hmc.edu/funfacts/ffiles/10006.3.shtml